perform a present worth analysis per the following project using a MARR of 8% The initial coast of the project is $100000, 60% of the initial cost will borrowed from a bonk using a loan that charges 5% interest componded annually. The loan will repaid with the equal annual payment .The project has life of 8 years. Revenne from the project will be $ 5000 in the first year and increasing by $3500 every year after that (ie: $5000 at the end of year 1,$8500 at the year 2, $12000 at Year 3) what the present worth of this project? Should this project be selected?
Answer #1

Solution:

Initial investment = $100,000 * 40% = $40,000

Bank Loan to finance the project = $100,000 * 60% = $60,000

Annual installment for loan repayment = $60,000 / Cumulative PV factor at 5% for 8 periods

= $60,000 / 6.463213 = $9,283

Computation of Present worth of Project
Particulars Total Year 1 Year 2 Year 3 Year 4 Year 5 Year 6 Year 7 Year 8
Annual Cash Inflows:
Revenue $5,000 $8,500 $12,000 $15,500 $19,000 $22,500 $26,000 $29,500
Repayment of bank loan -$9,283 -$9,283 -$9,283 -$9,283 -$9,283 -$9,283 -$9,283 -$9,283
Net Annual cash inflows -$4,283 -$783 $2,717 $6,217 $9,717 $13,217 $16,717 $20,217
PV Factor at 8% 0.925926 0.857339 0.793832 0.735030 0.680583 0.630170 0.583490 0.540269
Present value of cash inflows $37,708 -$3,966 -$671 $2,157 $4,570 $6,613 $8,329 $9,754 $10,923
Initial investment $40,000
Net Present worth -$2,292

As net present worth is negative, therefore project should not be selected.